Give an example, with justification, of a function with domain Z \{2,3}and codomain N.Is this function 1 –1?Is it onto ?Give reasons for your answers.


1
Expert's answer
2021-06-07T19:30:06-0400

Consider the function "f:\mathbb Z\setminus\{2,3\}\to\mathbb N,\ \ f(n)=\begin{cases} -n\ \text{ if }n< 0\\ n+1\ \text{ if }n\in\{ 0,1\} \text{ or } n\ge4 \end{cases}." By defenition of this function, the domain of "f" is "\mathbb Z\setminus\{2,3\}" and codomain is "\mathbb N."


Taking into account that "f(-1)=1" and "f(0)=1," we conclude that "f" is not 1-1. Since for any "n\in\mathbb N" we have that "f(-n)=-(-n)=n," the function "f" is onto.